OpenStudy (logan13):

What value of x makes the equation true? –7.12 = –4.8 + x A. –11.92 B. –11.2 C. –2.32 D. 2.32

OpenStudy (anonymous):

Add 4.8 to both sides.

OpenStudy (anonymous):

-7.12 + 4.8 = -4.8 + 4.8 + x

OpenStudy (anonymous):

Simplify both sides.

OpenStudy (chosenmatt):

eh i got here too late :(

OpenStudy (logan13):

um ya that never helped

OpenStudy (anonymous):

x = -7.12 + 4.8 -4.8 + 4.8 = 0, they cancel out..so this is what you're left with^

OpenStudy (anonymous):

Can you add -7.12 + 4.8?

OpenStudy (logan13):

no

OpenStudy (hockeychick23):

ok do 4.8-7.12

OpenStudy (logan13):

ʕ•́ᴥ•̀ʔっ

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (logan13):

no

OpenStudy (haseeb96):

MAy i help u @logan13

OpenStudy (anonymous):

Okay. -7.12 + 4.8 -12 + 8 = ?

OpenStudy (anonymous):

@ʕ•́ᴥ•̀ʔっ

OpenStudy (anonymous):

:D

OpenStudy (haseeb96):

-7.12 = -4.8+x when we will shift -4.8 on the other side its sign will be change so it would be -7.12 +4.8 =x x=-2.32 is your correct answer